mathslover (mathslover):

Prove that : \((n!)^2 > n^n\) for n>2

OpenStudy (anonymous):

you would have to use induction, correct?

mathslover (mathslover):

I don't think so. May be, induction may work here but I suppose, there is some other better method. Not sure!

OpenStudy (anonymous):

I see a problem with the statement. It is only true for n>2.

mathslover (mathslover):

Forgot to add that @Paynesdad .

OpenStudy (anonymous):

you could prove it by induction, that is the only way how i could see proving this inequality http://math.stackexchange.com/questions/170173/proving-inequalities-using-induction

OpenStudy (anonymous):

Would expressing the product as a sum help? Meaning, take the log of both sides, then you have the sum of \(n\) terms on the left and one term on the right. Something like \[2\sum_{k=0}^n\ln(n-k)>n\ln n\]

mathslover (mathslover):

What about using wavy curve method?

OpenStudy (anonymous):

Sorry, should be\[2\sum_{k=0}^{n-1}\ln(n-k)>n\ln n\] Not sure if it helps...

mathslover (mathslover):

Siths, yep, I am getting that but also not sure how to continue from that ...

OpenStudy (anonymous):

If anything, it might help in making a simpler induction proof. Starting with the base case \(n=3\), you have \[2\ln3+2\ln2+2\ln1>3\ln3\\ 2\ln2>\ln3\\ \ln4>\ln3\]

OpenStudy (anonymous):

If you can find a way of expressing this inequality for the \(n\)th term, I think you'll be in luck.

mathslover (mathslover):

Thinking...

OpenStudy (ankit042):

hmm I think I have a solution have a look first expand (n!)^2 as = n^2*(n-1)*[(n-2)*1]*[(n-2)*2]............[1*(n-2)] also write n^2 as n*n*n......n (n terms)

OpenStudy (amistre64):

hmm, since none ofthe parts equal zero\[(n!)^2 > n^n\] \[\frac{(n!)^2}{n^n} > 1\] \[\frac{n^2[(n-1)!]^2}{n^n} > 1\] \[n^{2-n}[(n-1)!]^2 > 1\] maybe?

OpenStudy (ankit042):

now first cancel n^2 from both the sides.....so on right we have n^n-2 Now divide n with each of the product pair i.e (n-2)*1/n.........(n-k-1)*(k-1)/n....1*(n-2)/n

mathslover (mathslover):

Nice @amistre64 and @ankit042 , I will respond about the solns you both have given, for now, i have to go for dinner :)

terenzreignz (terenzreignz):

Wait...rethinking...

OpenStudy (anonymous):

how about l'hopital?

OpenStudy (amistre64):

\[n^{2-n}[(n-1)!]^2 > 1\] \[[(n-1)!]^2 > \frac{1}{n^{2-n}}\] im sure this would be self evident :)

terenzreignz (terenzreignz):

It can't be that simple^ D:

OpenStudy (ankit042):

so on left we will have (n-1)[there will be n-2 pairs of type (n-k-1)k/n where 1<=k<=n-2] now the minimum value for a product pair will be when k =n/2 [THINK] so use this value for all k. We will get (n-1)(n-2)/2 >1 which is true for all n >2

OpenStudy (anonymous):

take the log, get \[2\ln(n!)>n\ln(n)\] or \[\frac{n\ln(n)}{2\ln(n!)}<1\] yeah that is silly, that is a limit true one though

OpenStudy (amistre64):

mines got a little flaw .. 2-n doesnt stay a fraction does it

OpenStudy (ankit042):

Guys this might not be very clear but have a look

OpenStudy (amistre64):

ahhh .... never try to solve math problems when you have to use the bathroom :)

terenzreignz (terenzreignz):

Okay, a new idea...Let \[\Large a_n = n^n\]\[\Large b_n = (n!)^2\] Then \[\Large \frac{a_n}{b_n}\] (after the dust settles) \[\Large \frac{a_n}{b_n}=\left( \frac{n+1}{n}\right)^n \cdot \frac1{n+1} \] Which, in terms of x... \[\Large \left( \frac{x+1}{x}\right)^x \cdot \frac1{x+1} \] Is a decreasing function from x \(\ge\) 2

terenzreignz (terenzreignz):

So, this \[\Large \frac{a_n}{b_n}\] is decreasing, as n goes to infinity, with n starting at 3. In fact, at n = 3, you have \[\Large \frac{a_n}{b_n}=\frac{27}{36}= \frac34 < 1\] And it's only going to get smaller as n gets larger, so, effectively, for all n \(\ge\) 3 \[\Large \frac{a_n}{b_n}<1\]\[\Large a_n <b_n\]\[\huge \color{blue}{n^n < (n!)^2}\]

mathslover (mathslover):

Nice one @terenzreignz